# Temperley-Lieb integrable models and fusion categories

### v2

@@ 4.1 Decomposition in terms of TL modules
We will furthermore restrict ourselves to even values of $L$, in which case $m$ takes integer values.
@@ 4.1.1 The Haagerup chain
We first discuss the TL chain built from the Haagerup $\mathcal{H}_{3}$ fusion category, discussed in section 2.4.2. The dimension of the corresponding Hilbert space scales as $(\frac{3+\sqrt{13}}{2})^{L}$. We obtain the decomposition of the Hilbert space into irreducible modules by comparing the Hamiltonian spectrum for finite sizes with the XXZ spectrum in the various irreducible modules, and recomposing the former in terms of the latter. For $L=4,6,8$, we find respectively
@@ 4.2 The low-energy spectrum
For real $\eta$ (equivalently, $\delta>2$, $\Delta>1$, which is the case for all three examples considered in this section), the XXZ chain is known to be gapped, and sits at the locus of a first order phase transition, which is that of the $Q$-states Potts model for $Q=\delta^{2}>4$ [56]. What this means for the energy spectrum is the following. In a sector of fixed magnetization and for a given twist, the lowest excited states are separated from the ground state by a gap which remains finite as $L\to\infty$. Similarly, changing the magnetization $m$ by a finite amount comes with a finite energy gap. However, within a given magnetization sector, changing the twist by a finite amount does not result in a gap. More quantitatively, one should expect that in the gapped phase the effect of a twist $\varphi$ (which, as discussed above, can be thought of as a local perturbation of magnitude $\varphi$ located on two sites of the chain) on the ground state energy decays as $e^{-L/\xi}$, where $\xi$ is the correlation length of the model, known from the exact solution [58] to be of the form
@@
in other words they can be quite large, and lead to important finite size effects when studying systems of, say, a few dozens of sites. In order to quantitatively observe the closure of the gap, we therefore use the underlying integrability of the models to compute the latter with Bethe ansatz.
@@ 4.3 Bethe ansatz study of the gap
In the Bethe ansatz, the eigenstates are constructed as interacting spin waves created on top of a reference state. They are characterised by a set of rapidities $\{\lambda_{1},\dots,\lambda_{N}\}$, where $N$ is related to the magnetization $m=S^{z}$ through N=L/2mN=L/2-m, while the rapidities satisfy the so-called Bethe equations
@@
The structure of the ground state for zero (as well as for real) twist is well-known [58, 63] : in the zero-magnetization sector, it is associated for real $\eta$ with a set of $N=L/2$ real rapidities distributed along the interval $\pi/2,\pi/2]$. For zero twist, those are symmetrically distributed with respect to the origin, while turning on a non-zero real twist results in an asymmetric distribution. Turning on an imaginary twist, this distribution deforms continuously and the rapidities becomes complex.
We can now readily use this technology to compute the energies $E_{0}(L)$ and $E_{1}(L)$ for large values of $L$. For a given value of $L$ and $\eta$, we start by computing the rapidities of the ground state at zero twist (the Bethe equations can be recast in an additive form involving a set of quantum numbers called “Bethe integers”, making their numerical resolution very easy [63]). From there, we vary the twist continuously, and obtain by a Newton-Raphson method the sets of rapidities for the ground states at twists $\varphi_{0}$ (imaginary) and $\varphi_{1}$ (real). Using the above expression for the energies, this gives access to the gap $\delta E(L)=E_{1}-E_{0}$. Since the expressions of $\varphi_{0}$ and $\varphi_{1}$ can be naturally extended to arbitrary $\eta$, we compute the latter for a range of values spanning the gapped phase of the XXZ model, and including the points $\eta_{\text{Fib}\times\text{Fib}}$ and $\eta_{\mathcal{H}_{3}}$. Our results are shown on on Figure 3, where for large enough $\Delta$ we clearly see the exponential vanishing of the gap,

### v3

@@ 4.1 Decomposition in terms of TL modules
We will furthermore restrict ourselves to even values of $L$, in which case $m$ takes integer values. We note that we obtained these decompositions into TL modules numerically on a case-by-case basis. Proving these mappings rigorously or finding a general pattern which holds for all fusion categories is an interesting open question for the future.
@@ 4.1.1 The Haagerup chain
We first discuss the TL chain built from the Haagerup $\mathcal{H}_{3}$ fusion category, discussed in section 2.4.2. The dimension of the corresponding Hilbert space scales as $(\frac{3+\sqrt{13}}{2})^{L}$. We obtain the decomposition of the Hilbert space into irreducible modules by comparing the Hamiltonian spectrum for finite sizes with the XXZ spectrum at different twist angles in the various irreducible modules, and recomposing the former in terms of the latter. For $L=4,6,8$, we find respectively
@@ 4.2 The low-energy spectrum
For real $\eta$ (equivalently, $\delta>2$, $\Delta>1$, which is the case for all three examples considered in this section), the XXZ chain is known to be gapped, and sits at the locus of a first order phase transition, which is that of the $Q$-states Potts model for $Q=\delta^{2}>4$ [56]. A qualitative understanding of the ground state and structure of low-energy excitations can be gained from perturbation theory at large coupling, $\Delta\to\infty$. To leading order the Hamiltonian is $H_{0}^{\rm XXZ}=\frac{\Delta}{2}\sum_{j}\sigma_{j}^{z}\sigma_{j+1}^{z}$, which has two degenerate ground states corresponding to the purely antiferromagnetic Néel and anti-Néel states. Excitations on top of the doubly degenerate ground state are built from domain walls between those two states, each domain wall coming at an energy cost $\Delta$. The next order contribution to the Hamiltonian flips pairs of adjacent spins, and vanishes on the ground state manifold. The ground states therefore remain degenerate to first order in $1/\Delta$, and separated by a finite gap from the rest of the system. Similar observations hold at higher orders, and in order to lift the ground state degeneracy, one has to go to order $1/\Delta^{L}$. This results in the following properties of the spectrum:
  • At a fixed value of the twist, the splitting between the two lowest energies is exponentially small in $L$, while higher energies are separated by a finite gap of order $\Delta$.
  • Changing the twist $\varphi$ by a finite amount results in an exponentially small shift in the ground state energy.
The same conclusions hold beyond perturbation theory, for any $\Delta>1$. More quantitatively, one should expect that in the gapped phase the effect of a twist $\varphi$ (which, as discussed above, can be thought of as a local perturbation of magnitude $\varphi$ located on two sites of the chain) on the ground state energy decays as $e^{-L/\xi}$, where $\xi$ is the correlation length of the model, known from the exact solution [58] to be of the form
@@
in other words they can be quite large, and lead to important finite size effects when studying systems of, say, a few dozen sites. In order to quantitatively observe the closure of the gap, we therefore use the underlying integrability of the models to compute the latter with Bethe ansatz.
@@ 4.3 Bethe ansatz study of the gap
In the Bethe ansatz, the eigenstates are constructed as interacting spin waves created on top of a reference state. They are characterised by a set of rapidities $\{\lambda_{1},\dots,\lambda_{N}\}$, where $N$ is related to the magnetization $m=S^{z}$ through N=L/2mN=L/2-m, while the rapidities satisfy the so-called Bethe equations [58, 63]
@@
The structure of the ground state and low-lying excitations of the XXZ chain at zero (as well as for real) twist is well-known [58, 63, 64]: in the zero-magnetization sector, the ground state is associated for real $\eta$ with a set of $N=L/2$ real rapidities distributed along the interval $\pi/2,\pi/2]$. These may be characterized by a set of $N$ distinct integer quantum numbers $I_{j}$, whose definition follows from taking the logarithm form of eqs. (51),
$L\phi_{\tfrac{\eta}{2}}(\lambda_{j})-\sum_{k\neq j}\phi_{\eta}(\lambda_{j}-\lambda_{k})=2\pi I_{j}+\varphi\,,\qquad\phi_{\chi}(\lambda)\equiv i\log\frac{\sin(\lambda+i\chi)}{\sin(\lambda-i\chi)}\,.$ (54)
For the ground state the integers $\{I_{j}\}$ are symmetrically packed around the origin, resulting for zero twist in a symmetric distribution of real Bethe roots, while turning on a non-zero real twist results in an asymmetric distribution. Turning on an imaginary twist, this distribution deforms continuously and the rapidities becomes complex. This is illustrated on Figure 3, where we show the root distributions associated with the ground state at $\varphi=0$ and $\varphi=i$ (blue and green dots respectively) for a given system size and anisotropy $\Delta$. We also display as yellow dots the Bethe roots associated with the exponentially degenerate ground state (“Umklapp” type excitation), obtained from the former by continuously driving the twist to $\varphi=2\pi$.
Figure 3: (Left) Low-lying spectrum of the XXZ chain obtained by exact diagonalization for $L=16$ sites, $\Delta=2.5$ and zero twist. The colored dots indicate the two exponentially degenerate ground states. (Right) The corresponding distributions of Bethe Roots in the complex plane. The blue and yellow dots correspond to the aforementioned ground states, while the green dots correspond to ground state root distribution for an imaginary twist $\varphi=i$.
Higher excited states, including those in sectors of nonzero magnetization, are obtained from the ground state by allowing for holes and strings in the distribution of Bethe roots, each coming with a finite energy cost [58, 63, 64] in agreement with the strong-coupling analysis presented in Section 4.2, together with the spectrum of Figure 3, obained by exact diagonalisation.
We can now readily use the Bethe ansatz technology to compute the energies $E_{0}(L)$ and $E_{1}(L)$ for large values of $L$. For a given value of $L$ and $\eta$, we start by computing the rapidities of the ground state at zero twist by numerically resolving the logarithmic Bethe equations (53) with a prescribed distribution of the integers $I_{j}$. From there, we vary the twist continuously, and obtain by a Newton-Raphson method the sets of rapidities for the ground states at twists $\varphi_{0}$ (imaginary) and $\varphi_{1}$ (real). Using the above expression for the energies, this gives access to the gap $\delta E(L)=E_{1}-E_{0}$. Since the expressions of $\varphi_{0}$ and $\varphi_{1}$ can be naturally extended to arbitrary $\eta$, we compute the latter for a range of values spanning the gapped phase of the XXZ model, and including the points $\eta_{\text{Fib}\times\text{Fib}}$ and $\eta_{\mathcal{H}_{3}}$. Our results are shown on on Figure 4, where for large enough $\Delta$ we clearly see the exponential vanishing of the gap,